OpenStudy (anonymous):

Together, teammates Pedro and Ricky got 2684 base hits last season. Pedro had 280 more hits than Ricky. How many hits did each player have?

OpenStudy (caozeyuan):

P+R=2684

OpenStudy (caozeyuan):

P-R=280

OpenStudy (caozeyuan):

2P=2684+280=2964

OpenStudy (caozeyuan):

P=1482

OpenStudy (caozeyuan):

R=1202

OpenStudy (anonymous):

Thankyou

OpenStudy (anonymous):

River C is 200 miles longer than River D. If the sum of their lengths is 5705 ​miles, what is the length of each​ river?

OpenStudy (caozeyuan):

this time you do it, use the same method as what I did, ok?

OpenStudy (caozeyuan):

Hint: start with setting up C=D and C-D

OpenStudy (caozeyuan):

sorry, C+D, not C=D

OpenStudy (anonymous):

ok thankyou

OpenStudy (anonymous):

the sum of their lengths is 5570 i messed up. but river c is 2885 miles right?

OpenStudy (caozeyuan):

sorry didnt see that, got it now or not?

OpenStudy (anonymous):

I just need help with river d

OpenStudy (caozeyuan):

if C is done, just subtract C from total